OpenStudy (anonymous):

The sum of the digits of a two-digit number is seven. When the digits are reversed, the original number is increased by nine. Find the number

OpenStudy (apoorvk):

okay let one's place digit be x and ten's place digit be y then the no. is 10y+x so, x+y=7 ....(i) after reversal the no. becomes 10x+y acc. to what's given, 10x+y=10y+x+9 simplifying, or, 9x-9y=9 or, x-y=1 ....(ii) solve equations (i) and (ii) to get x and y.

OpenStudy (apoorvk):

then ofcourse 10y+x is the no.!!

OpenStudy (anonymous):

34

OpenStudy (apoorvk):

Pikman. it helps if you dont just feed answers and can give an explanation :))

OpenStudy (anonymous):

if he needs an explanation then he would ask

OpenStudy (anonymous):

Yes true

OpenStudy (apoorvk):

ofcourse we arent just gonna feed answers. we give answers if asked not explnations! the concept is important my boy!! :)